Grade 12Organic Chemistry

A constant force of magnitude 50N is applied ona body of mass 10kg moving initially with a speed of 10 m s-1 against the direction of motion. how long will it take the body to come to rest?

raza
7 Years agoGrade 12

1 Answer

Yadesh
7 Years ago
Retardation produced in body is ==Force÷mass=50÷10=5m/s^2 
So now by eqn--
v=u+at
0=10+(--5)t
Or 10=5t
Hence time=2sec is answer